OpenStudy (alexandervonhumboldt2):

Guys i need help writting some math formulas on MathBB script:

OpenStudy (alexandervonhumboldt2):

here are the four i need to be written in MathBB script f1(x)=x^2 : R aroowtotheright R f2(x)=x^2 : R^+ aroowtotheright R f3(x)=x^2 : R aroowtotheright R^+ f4(x)=x^2 : R^+ aroowtotheright R^2

OpenStudy (alexandervonhumboldt2):

X right arrow Y

OpenStudy (alexandervonhumboldt2):

@Nnesha @inowalst @EclipsedStar @StudyGurl14

OpenStudy (alexandervonhumboldt2):

@poopsiedoodle

OpenStudy (poopsiedoodle):

well I have no idea how to mathbb so idk if I can help that much, but you could try the equation button there is this tho \(\rightarrow\) `\(\rightarrow\)`

OpenStudy (alexandervonhumboldt2):

thx

TheSmartOne (thesmartone):

\(\mathbb {f_1(x)=x^2 : R \rightarrow R}\) \(\mathbb{f_2(x)=x^2 : R^{+} \rightarrow R}\) \(\mathbb{f_3(x)=x^2 : R \rightarrow R^+}\) \(\mathbb{f_4(x)=x^2 : R^+ \rightarrow R^2}\) This might help you. And the codes for that is: ``` \(\mathbb {f_1(x)=x^2 : R \rightarrow R}\) \(\mathbb{f_2(x)=x^2 : R^{+} \rightarrow R}\) \(\mathbb{f_3(x)=x^2 : R \rightarrow R^+}\) \(\mathbb{f_4(x)=x^2 : R^+ \rightarrow R^2}\)

OpenStudy (unklerhaukus):

f4(x)=x^2 : R^+ aroowtotheright R^2 `f_4(x)=x^2 : \mathbb R^+ \to \mathbb R^2` \[\large f_4(x)=x^2 : \mathbb R^+ \to \mathbb R^2\]